what are number of ways of selcting 3 squares on a normal chess board so that theydont belong to same row column and diagonal line?explain

Grade:12

1 Answers

Saurabh Koranglekar
4 years ago
64C3-1148.

Calculate the ways in which 3 boxes can be selected which are in a straight line.
This can be divided into 2:
1) Diagonal=2*(3C3+4C3+5C3+....8C3)=252
2) Vertical/Horizontal
=2*(8*8c3)=896

Subtract this from the total number of ways of selecting 3 boxes on the chess board. This gives us the answer.